OpenStudy (anonymous):

integral of sin^5 3x cos 3x dx ??

OpenStudy (loser66):

let u = sin (3x) and step up from there

OpenStudy (anonymous):

3 cos 3x dx = du dx = 1 /3cos3x . dx it's true?

OpenStudy (loser66):

keep cos (3x) dx = du/3 and replace to integral

OpenStudy (anonymous):

the final answer is 1/6 sin^6 3x +c ??

OpenStudy (loser66):

\[\int \frac{u^5}{3}du=\frac{u^6}{6*3}+C = \frac{sin^6(3x)}{18}+C\]

OpenStudy (anonymous):

oh yeaah, thanks a lot :D

OpenStudy (loser66):

yw